calculate moles of 100g of electrons

Grade:12th pass

1 Answers

Harsh Patodia IIT Roorkee
9 years ago
1 e- is 9.1x10-31 kg

No of electrons in 100g i.e 0.1 kg = 0.1/(9.1x10-31)

6.022 x 1023 no of electron= 1 mole
.1/(9.1x10-31) electrons= ( 1/(6.022 x 1023) )x (.1/(9.1x10-31))
= 1.824x 105 moles